i do not know what to do to solve this example.. (root sign)9x(x to the second power) divided by (root sign) 18x(x to the second power)???
also there is this problem
2(rootsign)20 -3(rootsign)7 - 2(rootsign)5 + 4 (rootsign)63
ok i get 2(rootsign)5 - 3(rootsign)7 - 2(rootsign)5 + 3(rootsign)7 now does this become 2(rootsign)5 + 9(rootsign)7???

Answer by lotanez(23)

rewrite your problem:
√9x^2/√18x^2=(√9)(√x^2)/(√18)(√x^2)
=3x/3x√2
=3x(√2)/3x√2
=√2
√9=3
√x^2=x
√18=√9x2=3√2 (not a perfect square, so factor out the 18)
multiply both sides by √2 and the 3x cancel out.
2√20 - 3√7 - 2√5 + 4√63
= 2(√4)(√5) - 3√7 -2√5 + 4(√9)(√7) *√4=2, √9=3
= (2)(2)√5 -3√7 - 2√5 + (4)(3)√7
= 4√5 - 3√7 - 2√5 + 12√7
(you can only add the same radicals together)
= 4√5 - 2√5 - 3√7 + 12√7
= 2√5 + 9√7
Yes! You were right.